d yt) 2 dy() +7- + 10y() u(t) dt dt2 y(0) 0, y(0) = 3

Solve the following differential equation using variation of parameters.
Answer #1

The given difforental -egulion Solution (f)b = ROtBp t , y-3 dt dt2 2니 D&H0g= 4(8) cly-DR D+10)= 4f) (1) 2 auriliary eguation3 -201- S C3 57 201+503 --3- BTom equeHon (3) andl (), we 9et 20, + 2C2 O 20 +592 = -3 - 3 C 3 C2=-1 Purtting the val4e of c(f (FA P1Om eguatron (1 24 ult) e dt R)4() 31274 where u() is UhknownPutingr the value f u) 9nd u() in equation ), get 2ud 3 (dP SA R) dt 1 ( 3Then the g eneal 8olytion is 414) 40(4) +Ypl) -54 -24(24uldt -54 3 P Angwey